OpenStudy (anonymous):

What binomial is the factor of this trinomial ? x^2+5x+6

OpenStudy (math&ing001):

x^2+5x+6 = x^2+2x+3x+6 = x(x+2)+3(x+2) = (x+3)(x+2)